A single lane street 10 ft. wide goes through a semicircular tunnel with radius 9 ft. How high is the tunnel at the edge of each lane? Round off to 2 decimal places.
Answer by Alan3354(69443)

A single lane street 10 ft. wide goes through a semicircular tunnel with radius 9 ft. How high is the tunnel at the edge of each lane?
-----------
Each lane?
Says single lane.
----
The equation of the tunnel is x^2 + y^2 = 81
The height at each side is the y-value of the intersection of the circle and x = 5.
x^2 + y^2 = 81
25 + y^2 = 81
y^2 = 56
y = sqrt(56) =~ 7.48 feet